the length of a swimming pool is three times its wide less five meters,if it its volume is 304 cubic meters,its deep has 2 meters,determine the length and its wide. please, I need to find the solution using on Quadratic Formula.
please, answer to my e-mail : samuelpinedasarabia@gmail.com

Answer by mananth(16949)

let width be x m
length = 3x-5 m
Volume V = 304 cu.m
depth = 2m
...
Volume V = x(3x-5)*2 = 304
6x^2-10x=304
6x^2-10x-304 =0
/2
3x^2-5x-152 =0
...
a=3, b=-5,c=-152
..
x1= 5+sqrt (25+1824) / 2*3
x1=5+ sqrt(1849) / 6
x1=8
x2=-6.333 ( ignore this negative value)
width =8
length = 3*8-5
length = 19
...
C H E C K
8*19*2=304